OpenStudy (anonymous):

can someone please help me with this a) Calculate the final amount of a 20-year long-term savings plan if a payment of $2000 is made at the end of every year. Interest is 8.5% compounded quarterly. b) What happens to the final value of the long-term savings plan if a payment is made at the beginning of every year instead of at the end of every year?

OpenStudy (anonymous):

let's derive the formula by manually computing first years 1 st Year = D \[2\text{nd} \text{Year}= D\left(1+\frac{R}{4}\right)^4+D= D\left(\left(1+\frac{R}{4}\right)^4+1\right)\] \[3\text{rd} \text{Year}= D\left(\left(1+\frac{R}{4}\right)^4+1\right)\left(1+\frac{R}{4}\right)^4+D= D\left(\left(1+\frac{R}{4}\right)^{4*2}+\left(1+\frac{R}{4}\right)^4+1\right)\] \[4\text{th} \text{year}=\left(D\left(\left(1+\frac{R}{4}\right)^{4*2}+\left(1+\frac{R}{4}\right)^4+1\right)\right)\left(1+\frac{R}{4}\right)^4+D=D\left(\left(1+\frac{R}{4}\right)^{4*3}+\left(1+\frac{R}{4}\right)^{4*2}+\left(1+\frac{R}{4}\right)^4+1\right)\]

OpenStudy (anonymous):

\[\text{Nth} \text{year}=\text{ }D\left(1+\left(1+\frac{R}{4}\right)^4+\left(1+\frac{R}{4}\right)^{4*2}+\left(1+\frac{R}{4}\right)^{4*3}+\text{...}\text{...}+\left(1+\frac{R}{4}\right)^{4*(n-1)}\right)\] Rewritten using Geometric Sum Formula \[D\left(\frac{1-\left(\left(1+\frac{R}{4}\right)^4\right)^N}{1-\left(1+\frac{R}{4}\right)^4}\right)\]

OpenStudy (anonymous):

Plug in Numbers \[2000\left(\frac{1-\left(\left(1+\frac{.085}{4}\right)^4\right)^{20}}{1-\left(1+\frac{.085}{4}\right)^4}\right)\]=99770.2

OpenStudy (dumbcow):

yes this is correct

OpenStudy (dumbcow):

if it was made at beginning of each year, the difference in the formula would be that the first term in the sequence would be (1+.085/4)^4 not 1 and each consecutive terms exponent would be shifted up 1 =108,524.9

OpenStudy (dumbcow):

if you are able, financial calculators are useful for these type of problems to check your work...i use a TI BAII

OpenStudy (anonymous):

thanks buddy